SPECTROSCOPIC CHARACTERIZATION OF Be_{2}^{+} X^{2}Σ_{u}^{+} AND THE IONIZATION ENERGY OF Be_{2}

\underline{I. O. ANTONOV}, B. J. BARKER, V. E. BONDYBEY, M. C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Rotationally resolved spectra for Be_{2}^{+} were recorded with PFI-ZEKE technique. Vibrational levels v^{+}=0-6 were observed. The symmetry of the ground state was determined as ^{2}Σ_{u}^{+}. The bond energy was found to be D_{e}^{+} = 16348(5) cm^{−1} and the equilibrium distance R_{e}^{+} = 2.211(8) Å. The ionization energy for Be_{2} was refined at 59824(2) cm^{−1}. Comparisons with high-level theoretical calculations indicate that the bonding in Be_{2}^{+} is adequately described by MRDCI calculations with moderately large basis sets.